OpenStudy (anonymous):

solve the system by graphing -2x=2y-4 2x-y=-5 *(-1,3) *(3,-1) *(-3,-1) *(-1,-3)

OpenStudy (solomonzelman):

graph them and at the point where they intersect is the answer. to graph them, put them in y=mx+b from and plug in x values.

OpenStudy (ivettef365):

try the points, substitute x and y to see which one is true

OpenStudy (anonymous):

i think its B, is that right?

OpenStudy (ivettef365):

lets try B (3,-1) (x,y) -2x=2y-4 = -2(3) = 2(-1) -4 -6 = -2 - 4 -6 = -6 ok, works 2x-y=-5 = 2(3) -(-1) = -5 6 + 1 = -5 7 = -5 not good need to keep trying

OpenStudy (anonymous):

ok